### v1

# Gravitational and electromagnetic Cherenkov radiation constraints in modified dispersion relations

Abstract
Motivated by different approaches to quantum gravity, one could consider that Lorentz invariance is not an exact symmetry of nature at all energy scales. Following this spirit, modified dispersion relations have been used to encapsulate quantum gravity phenomenology. In the present work, we propose a class of Lorentz invariance violating phenomenological dispersion relations, which could be different for each particle species, to study the generalized vacuum Cherenkov radiation process. We identify the kinematic regions where the process is allowed and then compute the energy loss rate due to the emission of vacuum electromagnetic and gravitational Cherenkov radiation. Furthermore, we estimate constraints for the Lorentz invariance breaking parameters of protons and gravitons taking into account the existence (or absence) of vacuum gravitational Cherenkov radiation using ultra high energy cosmic ray detections.
preprint: IPARCOS-UCM-24-044
1 Introduction
Electromagnetic Cherenkov radiation occurs when an electrically charged particle travelling through an optical medium has a velocity $v$ larger than the phase velocity of light in that medium $c_{\gamma}$. This phenomenon was first detected by Cherenkov in 1934 Cherenkov_1934 , and a theoretical explanation was given by Frank and Tamm three years later Tamm_1937 . Cherenkov radiation (CR) can be understood microscopically in the following way Schreck_2018 . When the particle travels at a velocity $v\leq c_{\gamma}$, the polarized atoms close to the trajectory emit out of phase radiation, leading to destructive interference. However, when the particle has a velocity $v>c_{\gamma}$, the wave trains emitted by atoms and molecules are in phase, and the constructive interference taking place in a Mach cone, known as the Cherenkov cone, results in the coherent radiation observed in the direction perpendicular to the cone. Likewise, this phenomenon can be interpreted both qualitatively and quantitatively as an emission process $a\to a+\gamma$ where a charged particle $a$ emits a photon $\gamma$: the particle outruns the electromagnetic field, causing the emission of radiation because of the accumulation of wavefronts propagating from the particle Schreck_2018 ; Moore_2001 . The rate of energy loss in an optical medium with refractive index $n$ is described classically by the Frank-Tamm formula Tamm_1937 , and receives small corrections when considering quantum effects cox1944momentum . Moreover, this formula can also be obtained from the quantum field theory formalism Moore_2001 . Anyway, for the process to be allowed the velocity of the particle must be greater than the phase velocity of light and, therefore, Cherenkov emission is not possible in vacuum in the framework of special relativity (SR) since $c_{\gamma}=c$.
Nevertheless, over the last decades there has been a growing interest in studying theoretical frameworks which suggest that Lorentz invariance (LI) may not be an exact symmetry of nature at all energy scales Mattingly_2005 . Such theories appear mainly in the context of quantum gravity (QG) Addazi_2022 ; batista2023white where, if LI is violated, the Planck energy $E_{\text{Pl}}\approx 10^{19}\,\text{GeV}$ is expected to be the scale where this symmetry is strongly violated. Although there is a large energy gap between the highest energy particles detected, those are Ultra High Energy Cosmic Rays (UHECRs) $\approx 10^{11}\,\text{GeV}$, and the Planck scale, there should be an interpolation of LI violation (LIV) to the low energy regime, where these particles could be sensitive to small departures from LI. Focusing on astrophysical phenomena, the most straightforward way to implement LIV is to consider modified dispersion relations (MDRs) for particles maintaining the $4-$momentum conservation laws. In this scenario, kinematics of a wide variety of processes may be affected in what are called threshold effects depending on the MDRs considered. These effects could entail the shift of the energy thresholds or add new thresholds to processes already predicted in the LI scenario, or even allow completely new reactions.
Vacuum CR may be allowed in a LI violating frame, and it has been exhaustively studied in the electromagnetic sector Coleman_1997 ; Coleman_1999 ; Jacobson_2003 ; Jacobson_2006 ; Gagnon_2004 to impose restrictive constraints to LI violating parameters (LIVPs) using high energy astrophysical particles. This idea has also been extended to study the vacuum gravitational CR Moore_2001 ; Elliott_2005 ; Kimura_2012 ; Kiyota_2015 ; Kosteleck_2015 , where an arbitrary particle can lose energy due to the emission of gravitons rather than photons. With the recent observations of gravitational waves (GWs) Yunes_2016 ; Abbott_2017 , combined constraints for LIVPs were obtained in the gravitational sector. These studies, however, only modify the dispersion relation in the gravitational sector or the matter sector, but combined effects have not been investigated yet with non-trivial MDRs. The main aim of the present paper is to consider MDRs both in the gravitational and the matter sector simultaneously to obtain combined restrictions in the LIVPs taking into account the existence or absence of vacuum gravitational CR.
The remainder of this work is structured as follows. In section 2 we propose a phenomenological MDR and study the kinematics of the generalised Cherenkov radiation (GCR) process $a\to a+b$. In section 3 we apply the results obtained in section 2 to compute the rate of energy loss of the particle (section 3.1) for vacuum electromagnetic and gravitational CR (sections 3.2 and 3.3). Section 4 is devoted to impose constraints in LIVPs using UHECRs, and, finally, in section 5 we discuss the main results obtained and its limitations. Along this work we use natural units, i.e., we set $\hbar=c=1$.
2 Kinematics of generalised Cherenkov radiation
2.1 General aspects of the modified dispersion relation
Nowadays it is not yet clear how LI might be broken, if it is the case. However, there are hints about the kind of phenomenology it could produce. An interesting example is the possibility of new particle decays due to LIV. In order to study the kinematics of these processes, it has previously been proposed Coleman_1999 that each particle $a$ has, besides its own mass $m_{a}$, its own maximum velocity $c_{a}$, which is asymptotically achievable if the particle has a non-vanishing mass. The dispersion relation in this scenario is given by
$E_{a}^{2}=m_{a}^{2}c_{a}^{4}+\bm{p}_{a}^{2}c_{a}^{2},$ (1)
where $E_{a}$ and $\bm{p}_{a}$ are the energy and 3-momentum of the particle, respectively. On the other hand, the Cherenkov effect is usually studied in terms of the refractive index $n$ of an optical medium, which modifies the maximum attainable velocity of light in vacuum $c$ as $c_{\gamma}=cn^{-1}$. Hereafter we will set $c=1$. Following this idea, we consider that it is convenient to parameterize the deviation of the maximum attainable velocity $c_{a}$ for each particle species in terms of a particular refractive index $n_{a}$, so that $c_{a}=n_{a}^{-1}$ and equation (1) can be written as
$E_{a}^{2}=m_{a}^{2}n_{a}^{-4}+\bm{p}_{a}^{2}n_{a}^{-2}.$ (2)
The existence of this refractive index $n_{a}$ for particles of the Standard Model can be interpreted in the context of quantum gravity, for instance, as the spacetime foam acting as a medium Ellis_2008 . In a similar manner, for GWs it may act as a diagravitational medium Cembranos_2019 .
In this work we propose a phenomenological dispersion relation where the refractive index $n_{a}$ of a particle $a$ can depend on its energy as a power law
$n_{a}=1+\mathbb{A}^{(\alpha)}E_{\bm{p}_{a}}^{\alpha},$ (3)
where there is no summation in $\alpha$ as we take into account only the dominant term in the deviation from unity. Both the coefficient $\mathbb{A}^{(\alpha)}$, with units $[\mathbb{A}^{(\alpha)}]=[E^{-\alpha}]$, and the exponent $\alpha$ are allowed to be different for each particle species. From now on we shall denote $\mathbb{A}\equiv\mathbb{A}^{(\alpha)}$ except when discussing different values of $\alpha$. As LI is almost an exact symmetry we will assume that $|\mathbb{A}E_{\bm{p}_{a}}^{\alpha}|\ll 1$. Replacing the refractive index (3) in the MDR (2) and considering terms up to first order in $|\mathbb{A}E_{\bm{p}_{a}}^{\alpha}|$ we get
$E_{\bm{p}_{a}}^{2}=m_{a}^{2}+\bm{p}_{a}^{2}-2\mathbb{A}E_{\bm{p}_{a}}^{\alpha}(2m_{a}^{2}+\bm{p}_{a}^{2}).$ (4)
Even though $m_{a}$ and $|\mathbb{A}E_{\bm{p}_{a}}^{\alpha}|$ are independent scales, we will see that it is safe to expand only in the LI violating term for high energy particles, hence no multiscale problems arise from this analysis. Note that equation (4) is invariant under translations and rotations but not under Lorentz boosts due to the dependence of $n_{a}$ on the energy of the particle. Thus, this MDR only holds in a particular preferred frame, which is usually chosen to be in rest with respect to the CMB. The relative velocity between Earth and the CMB is of order $\mathcal{O}(10^{-3})$ Planck_2018 and its effects can be neglected when imposing constraints on the LIVPs.
It should be noted that for the MDR given by equation (2) the group velocity and the phase velocity are different in general. The maximum attainable velocity of the particle $c_{a}=n_{a}^{-1}=1-\mathbb{A}p^{\alpha}+\mathcal{O}(\mathbb{A}^{2})$ is referred to the phase velocity, which differs from the group velocity defined as $v_{a}(E)=\partial E/\partial p$. Indeed, for high energy particles with $p\gg m$ one has, using equation (4) and neglecting the mass in the last term,
$v_{a}=1-\frac{m^{2}}{2p^{2}}-(\alpha+1)\mathbb{A}p^{\alpha}+\mathcal{O}(\mathbb{A}^{2}),$ (5)
where $p$ is the module of the $3-$momentum of the particle. We then see that the maximum achievable group velocity, $v_{a}(m=0)=1-(\alpha+1)\mathbb{A}\,p^{\alpha}$, is different from that given by $c_{a}$ when $\alpha\neq 0$. As we will see, for the processes we want to study, the kinematics is governed by the relation between the maximum attainable phase velocities $c_{a}$ of the particles, or, equivalently, the relation between the refractive indexes $n_{a}$.
The MDR (4) could be used to study the shift in the threshold energy or new thresholds of existing reactions. If the mass of the particle is high enough compared to the LI violating term, then the usual dispersion relation holds and kinematics remain, essentially, unmodified. However, if the LI breaking term is comparable to the mass of the particle, significant deviations are expected in the kinematics of a wide variety of reactions. As the thresholds of the processes are determined by the mass of the particles, modifications to those thresholds are appreciable when the third term in the MDR (4) is of the same order as the mass $m_{a}$. For high energy particles $p\gg m$ the energy $p_{\text{dev}}$ at which the deviation becomes important can be estimated by
$p_{\text{dev}}\sim\left|\frac{m^{2}}{2\mathbb{A}}\right|^{1/(\alpha+2)}.$ (6)
In this work we are interested in studying the existence of entirely new reactions that are forbidden in SR; in particular we will consider the GCR process (see figure 1), i.e., the two body decay process $a\to a+b$ where a particle $a$ emits a massless particle $b$ [^1] Here we mention that forbidden processes in the LI scenario cannot be studied in the scope of Doubly Special Relativity (DSR). This is because the deformed $4-$momentum composition produces cancellations with the LI violating terms in such a way that this phenomenology is not possible Addazi_2022 ..
Figure 1: Tree level GCR process where a particle $a$ (continuous line) emits a massless particle $b$ (wiggly line).
2.2 Derivation of the Cherenkov angle
One of the key ingredients when studying CR is the angle of emission $\theta_{c}$ between the $3-$momentum $\bm{p}$ of the initial particle $a$ and the $3-$momentum $\bm{k}$ of the emitted particle $b$. If we define $E_{\bm{p}}:=E_{a}(\bm{p})$, $E_{\bm{p^{\prime}}}:=E_{a}(\bm{p}^{\prime})$ and $E_{\bm{k}}:=E_{b}(\bm{k})$, where $\bm{p}^{\prime}$ is the $3-$momentum of the particle $a$ after the emission (see figure 1), then the conservation of energy and $3-$momentum, which are written as
$E_{a}(\bm{p})=E_{b}(\bm{k})+E_{a}(\bm{p}^{\prime}),\hskip 15.0pt\bm{p}=\bm{k}+\bm{p}^{\prime},$ (7)
allow us to obtain $\theta_{c}$ in terms of the module of the $3$-momenta $\bm{p}$ and $\bm{k}$. Note that equation (7) assumes the usual $4-$momentum composition, hence DSR theories are excluded from this analysis.
Let us now use the MDR given by equation (2) with refractive index (3) to obtain the Cherenkov angle $\theta_{c}$ or, equivalently, $\cos{\theta_{c}}$ ). Let $m$ be the mass of particle $a$ (we remind the reader that $b$ is massless) and denote $n_{\bm{p}}:=n_{a}(\bm{p})$, $n_{\bm{p}-\bm{k}}:=n_{a}(\bm{p}-\bm{k})$ and $n_{\bm{k}}:=n_{b}(\bm{k})$, where we already applied the conservation of $3-$momentum. From equation (2), we see that the energies of the particles in the GCR process are given by
$E_{\bm{p}}^{2}$ $=m^{2}n_{\bm{p}}^{-4}+p^{2}n_{\bm{p}}^{-2},$
$E_{\bm{p}-\bm{k}}^{2}$ $=m^{2}n_{\bm{p}-\bm{k}}^{-4}+(p^{2}+k^{2}-2pk\cos{\theta_{c}})n_{\bm{p}-\bm{k}}^{-2},$ (8)
$E_{\bm{k}}^{2}$ $=k^{2}n_{\bm{k}}^{-2},$
where we have used $(\bm{p}-\bm{k})^{2}=p^{2}+k^{2}-2pk\cos{\theta_{c}}$ ) and denoted the module of the $3-$momentum of the particles $a$ and $b$ as $p$ and $k$, respectively. On the other hand, combining the conservation of energy and $3-$momentum we have $E_{\bm{p}-\bm{k}}^{2}=[E_{\bm{p}}-E_{\bm{k}}]^{2}$. Substituting here the energies given by equation (8), we get
$2pkn_{\bm{p}-\bm{k}}^{-2}\cos{\theta_{c}}$ ) $=$ $p^{2}(n_{\bm{p}-\bm{k}}^{-2}-n_{\bm{p}}^{-2})+m^{2}(n_{\bm{p}-\bm{k}}^{-4}-n_{\bm{p}}^{-4})+k^{2}(n_{\bm{p}-\bm{k}}^{-2}-n_{\bm{k}}^{-2})$ (9)
$+$ $2n_{\bm{k}}^{-1}k\left[n_{\bm{p}}^{-2}p^{2}+n_{\bm{p}}^{-4}m^{2}\right]^{1/2}.$
Note that in the LI scenario, which corresponds to $n_{a}=n_{b}=1$, the RHS of equation (9) is always greater than the LHS unless $\cos{\theta_{c}}>1$; hence, the process is forbidden. The limit case where the particle $a$ has no mass is of no interest when $n_{a}=n_{b}=1$, as the emission rate, which we compute in section 3, vanishes for $\cos{\theta_{c}}=1$.
Now let $\mathbb{A}$ and $\alpha$ be the LIVP and exponent of the refractive index of the particle $a$, and $\mathbb{B}$ and $\beta$ those of the particle $b$, such that
$n_{a}(\bm{p})=1+\mathbb{A}E_{a}^{\alpha}(\bm{p}),\hskip 15.0ptn_{b}(\bm{k})=1+\mathbb{B}E_{b}^{\beta}(\bm{k}).$ (10)
We remark that for the massless particle we are ignoring a small mass term $m_{b}$, and terms that induce birefringence or attenuation, which could be present in alternative theories of gravity for the graviton Cembranos_2019 or in the Standard Model extension for gauge bosons. Assuming that $|\mathbb{A}E_{\bm{p}}^{\alpha}|,|\mathbb{A}E_{\bm{p}-\bm{k}}^{\alpha}|,|\mathbb{B}E_{\bm{k}}^{\alpha}|\ll 1$ we can replace these refractive indexes in equation (9) and expand up to first order to obtain
$2pk\cos{\theta}$ ) $=$ $4pk\mathbb{A}E_{\bm{p}-\bm{k}}^{\alpha}\cos{\theta}-2\mathbb{A}(p^{2}+2m^{2})[E_{\bm{p}-\bm{k}}^{\alpha}-E_{\bm{p}}^{\alpha}]$ (11)
$+$ $2k(p^{2}+m^{2})^{1/2}\left[1-\mathbb{B}E_{\bm{k}}^{\beta}\right]\left[1-2\mathbb{A}\frac{(p^{2}+2m^{2})E_{\bm{p}}^{\alpha}}{p^{2}+m^{2}}\right]^{1/2}$
-- $2k^{2}[\mathbb{A}E_{\bm{p}-\bm{k}}^{\alpha}-\mathbb{B}E_{\bm{k}}^{\beta}].$
This is an implicit equation in $\cos{\theta}$ ) as the energy $E_{\bm{p}-\bm{k}}$ depends on $(\bm{p}-\bm{k})^{2}$, which in turn depends on $\cos{\theta}$ ).
In section 4 we will consider UHECR observations to constrain the LI breaking parameters $\mathbb{A}$ and $\mathbb{B}$. For these particles $p\gg m$. In addition, it can be checked a posteriori that the threshold is greatly modified, that is, the momentum is much larger than the right hand side of equation (6). We can then consider terms up to $\mathcal{O}(m^{2}/p^{2})$ and neglect first order products of the LI violating terms with the quotient $m^{2}/p^{2}$. Applying these additional approximations in equation (11) we get
$\cos{\theta_{c}}:=1-\Theta_{c}$ (12)
where
$\Theta_{c}:=\frac{p-k}{pk}\left\{(p-k)\mathbb{A}E_{\bm{p}-\bm{k}}^{\alpha}+\mathbb{B}kE_{\bm{k}}^{\beta}-\mathbb{A}pE_{\bm{p}}^{\alpha}\right\}-\frac{m^{2}}{2p^{2}}.$ (13)
As $\Theta_{c}$ is already of first order in the LI violating terms and $m^{2}/p^{2}$, then the $3-$vectors $\bm{p}$ and $\bm{k}$ involved in the energies must be considered collinear and no second order corrections are taken into account. In other words, $\Theta_{c}$ does not depend on $\cos{\theta_{c}}$ ) and, consequently, equation (12) completely characterizes the Cherenkov angle $\theta_{c}$ in terms of the initial $3-$momentum of the particle $a$ and the $3-$momentum carried off by the particle $b$.
2.3 Threshold condition
Attending to equation (12), the Cherenkov process $a\to a+b$ is allowed as long as $\Theta_{c}>0$ such that $\cos{\theta_{c}}<1$. Hence the threshold is determined by the equation
$\Theta_{c}=0.$ (14)
The study of the threshold condition is easier to carry out if the energies in equation (13) are substituted by the momentum, which is correct up to first order. Thus, we obtain
$\Theta_{c}$ $=$ $\frac{p-k}{p}f(k)-\frac{m^{2}}{2p^{2}},$ (15)
$f(k)$ ::=:= $\mathbb{B}k^{\beta}-\frac{\mathbb{A}}{k}\left\{p^{\alpha+1}-(p-k)^{\alpha+1}\right\}.$ (16)
In what follows we will consider $k\leq p$, which is satisfied particularly when studying the threshold condition. Indeed, if $k>p$, then the massless particle would be anti-parallel to the incoming particle, but this is forbidden by the threshold theorem Mattingly_2003 : if $E_{\bm{p}}$ is a strictly monotonically increasing function of $p$ for $p>0$ for all particles, then all thresholds for processes with two particle final states occur when the final momentum are parallel. This theorem has to be satisfied in our case since, in the first place, as we have previously discussed, $E_{\bm{p}}$ is a rotational-invariant function of $\bm{p}$, so it is only function of the modulus of the 3-momentum $p$; in the second place, it is a strictly monotonically increasing function of $p$, as the LI violating terms are much smaller than the momentum of the particle.
Before focusing on the threshold condition, it is possible to establish whether the process is allowed or forbidden attending to the signs of the parameters $\mathbb{A}$ and $\mathbb{B}$. Two cases can be studied without loss of generality.
  1. 1.
    $\mathbb{A}>0$ and $\mathbb{B}<0$. It is easy to see from equation (15) that $\Theta_{c}<0$ for all $k<p$, as the energy of the incoming particle is always greater than the energy of the particles in the final state. Hence, Cherenkov emission is forbidden. In this scenario $n_{a}>n_{b}$ and, therefore, $c_{a}<c_{b}$. So, the intuitive condition that the particle $a$ must have a greater phase velocity than the particle $b$ in order for the process to be allowed is never fulfilled.
  2. 2.
    $\mathbb{A}<0$ and $\mathbb{B}>0$. In this case we see that all the terms in equation (15) are positive except the correction $m^{2}/(2p^{2})$ due to the mass of the particle. For particles with energies many orders of magnitude greater than the one established by the RHS of equation (6), this term might be neglected and, thus, Cherenkov emission is allowed for $k\lesssim p$. Now we have that $n_{a}<n_{b}$ ($c_{a}>c_{b}$), which fits with the idea that the process is possible since the particle $a$ has a greater phase velocity than the particle $b$.
[…]
Before doing so, we can study the relation between $\mathbb{A}$ and $\mathbb{B}$ when the emitted particle $b$ has arbitrarily low momentum, $k\to 0$, or carries off all of the momentum of the particle $a$, $k\to p$. In the former case, it is easy to see from equation (15) that Cherenkov emission is possible if:
[…]
Note that the first case corresponds to a constant refractive index $n_{b}$. The process is not allowed when $k=p$ and $m\neq 0$, but for sufficiently high energy particles such that the mass may be neglected the process is allowed as long as
$\mathbb{B}p^{\beta}-\mathbb{A}p^{\alpha}\geq 0.$ (18)
The threshold condition is recovered setting an equal sign in equations (17) and (18).
It is also possible to prove in the massless limit that if the threshold condition $f(k_{\text{th}})=0$ has a solution, then $k_{\text{th}}$ is unique. Let us first focus on the case where $\mathbb{A}>0$ and $\mathbb{B}>0$. We start by calculating the derivative of $f(k)$, which is given by
$f^{\prime}(k)$ $=\beta\mathbb{B}k^{\beta-1}+\mathbb{A}k^{-2}p^{\alpha+1}g(k),$ (19)
$g(k)$ $=1-\left(1-\frac{k}{p}\right)^{\alpha}\left(1+\frac{\alpha k}{p}\right).$ (20)
It is easy to check that, for the physical momentum $k\in[0,p]$, $g^{\prime}(k)>0$ and $g(0)$, $g(p)>0$; hence $g(k)>0$. The remaining terms in equation (19) are positive, so we also have that $f^{\prime}(k)>0$; in other words, $f(k)$ is a monotonically increasing function. Therefore, if $f(0)<0$ and $f(p)>0$, the solution $k_{\text{th}}$ is unique. The condition $f(0)<0$ is satisfied when equation (17) does not hold, and $f(p)>0$ when equation (18) is fulfilled. The threshold momentum $k_{\text{th}}$ here establishes the minimum momentum that particle $b$ must have in order to have Cherenkov emission with momentum $k\in[k_{\text{th}},p]$, and it is permitted as long as $k_{\text{th}}<p$. The limit case $k_{\text{th}}=p$ occurs setting an equal sign in equation (18).
The prove is similar when $\mathbb{A}<0$ and $\mathbb{B}<0$. In this case one can check that $f(k)$ is a monotonically decreasing function, and so there exists a unique solution $k_{\text{th}}$ when $f(0)>0$ and $f(p)<0$. The first condition is met when equation (17) is satisfied, and the second one when equation (18) is not fulfilled. Hence, $k_{\text{th}}$ establishes the maximum momentum that particle $b$ can have up to where the Cherenkov effect is allowed, and thus $k\in[0,k_{\text{th}}]$. The process is permitted when $k_{\text{th}}>0$, and the limit case $k_{\text{th}}=0$ is given setting an equal sign in equation (17).
When mass is considered in the previous analysis, one has to take into account that $\Theta_{c}(p)<0$ always. The discussion when $\mathbb{A}<0$ and $\mathbb{B}<0$ holds: if $f(0)>0$ there is only one solution to the thresholds condition and it is the maximum momentum that particle $b$ can have. In the case where $\mathbb{A}>0$ and $\mathbb{B}>0$, when $f(0)<0$ it is possible to have two solutions, so that $k\in[k_{\text{min}},k_{\text{max}}]$ with $k_{\text{max}}<p$, or no solutions.
$\alpha$ $\beta$ $k_{\text{th}}$ $\mathbb{A},\mathbb{B}>0$ $\mathbb{A},\mathbb{B}<0$
00 00 No threshold $\mathbb{B}-\mathbb{A}>0$ $\mathbb{B}-\mathbb{A}>0$
00 $\mathbb{R}^{+}$ $\left(\frac{\mathbb{A}}{\mathbb{B}}\right)^{1/\beta}$ $\mathbb{B}p^{\beta}-\mathbb{A}>0$ $\mathbb{A}<-\frac{m^{2}}{2p^{2}}$
$1$ 00 $2p-\frac{\mathbb{B}}{\mathbb{A}}$ $\mathbb{B}-\mathbb{A}p>0$ $\mathbb{B}-2\mathbb{A}p<\frac{m^{2}}{2p^{2}}$
$1$ $1$ $\frac{2\mathbb{A}p}{\mathbb{A}+\mathbb{B}}$ $\mathbb{B}p-\mathbb{A}p>0$ $\mathbb{A}<-\frac{m^{2}}{4p^{3}}$
$1$ $2$ $\frac{1}{2\mathbb{B}}\left(-\mathbb{A}\pm\sqrt{\mathbb{A}^{2}+8\mathbb{A}\mathbb{B}}\right)$ $\mathbb{B}p^{2}-\mathbb{A}p>0$ $\mathbb{A}<-\frac{m^{2}}{4p^{3}}$
$2$ 00 $\frac{1}{2\mathbb{A}}\left(3\mathbb{A}p\pm\sqrt{4\mathbb{A}\mathbb{B}-3\mathbb{A}^{2}p^{2}}\right)$ $\mathbb{B}-\mathbb{A}p^{2}>0$ $\mathbb{B}-3\mathbb{A}p^{2}>\frac{m^{2}}{2p^{2}}$
$2$ $1$ $\frac{1}{2\mathbb{A}}\left(3\mathbb{A}p+\mathbb{B}\pm\sqrt{\mathbb{B}^{2}+6\mathbb{A}\mathbb{B}p-3\mathbb{A}^{2}p^{2}}\right)$ $\mathbb{B}p-\mathbb{A}p^{2}>0$ $\mathbb{A}<-\frac{m^{2}}{6p^{4}}$
$2$ $2$ $\frac{1}{2(\mathbb{A}-\mathbb{B})}\left(3\mathbb{A}p\pm\sqrt{12\mathbb{A}\mathbb{B}-3\mathbb{A}^{2}p^{2}}\right)$ $\mathbb{B}p^{2}-\mathbb{A}p^{2}>0$ $\mathbb{A}<-\frac{m^{2}}{6p^{4}}$
Table 1: In this table we show the threshold momentum $k_{\text{th}}$ for the particle $b$ that satisfy the condition $f(k)=0$ together with the conditions that $\mathbb{A}$ and $\mathbb{B}$ should satisfy for the process to be allowed. The $\pm$ signs in the quadratic solutions depend on the signs of $\mathbb{A}$ and $\mathbb{B}$, and these are chosen so that $k_{\text{th}}$ is positive for those LIVPs that fulfill either the condition (17) or (18). We emphasize again that $k_{\text{th}}$ establishes the minimum momentum of the particle $b$ when $\mathbb{A},\mathbb{B}>0$, and a maximum momentum when $\mathbb{A},\mathbb{B}<0$.
The threshold condition $\Theta_{c}=0$ does not have a general solution $k_{\text{th}}$ for arbitrary $\alpha$ and $\beta$ since it is a polynomial equation of degree $\max\{\alpha+1,\beta+1\}$. Nevertheless, particular solutions for small values of $\alpha$ and $\beta$ can be derived, and further simplified in the massless limit $m\to 0$. For example, consider the case where $\alpha=1$ and $\beta=0$. One easily checks from equation (16) that the threshold momentum for $m=0$ is given by
$k_{\text{th}}(\alpha=1,\beta=0)=2p-\frac{\mathbb{B}}{\mathbb{A}}.$ (21)
Let us start discussing the case $\mathbb{A},\mathbb{B}>0$, where the threshold momentum is the minimum value of $k$ for the process to be allowed. Then, $k_{\text{min}}>p$ if $\mathbb{B}<\mathbb{A}p$ and the process is forbidden, as can be seen from equation (18). Otherwise, the process is allowed with minimum momentum given by equation (21); in particular, the process is allowed for all $k$ when $\mathbb{B}>2\mathbb{A}p$, see equation (17). The analysis is reversed when considering $\mathbb{A},\mathbb{B}<0$. Now equation (21) establishes the maximum momentum up to where the process is allowed. As long as $\mathbb{B}<2\mathbb{A}p$ the process is permitted, and for $\mathbb{B}<\mathbb{A}p$ it is possible for all values of $k$; for $\mathbb{B}>2\mathbb{A}p$ Cherenkov emission is not possible as $k_{\text{max}}<0$. A similar analysis can be carried out solving equation (14) for different values of $\alpha$ and $\beta$, and study whether the process is allowed or not attending to the signs of $\mathbb{A}$ and $\mathbb{B}$ and to equations (17) and (18). Table 1 illustrates the solutions to the threshold condition in the massless limit for the values $\alpha,\beta\in\{0,1,2\}$.
3 Dynamics of generalised Cherekov radiation
In this section we focus on the calculation of the rate of energy loss $\text{d}E/\text{d}t$ of the particle $a$ in the GCR process $a\to a+b$. We present the general procedure for the tree-level diagram shown in figure 1 for an arbitrary interaction vertex. Then we study the particular cases where the emitted particle is either a photon $\gamma$ (electromagnetic CR) or a graviton $h$ (gravitational CR).
3.1 Decay rate and energy loss
It is well known peskin1995introduction that the differential decay rate of a two body process in an arbitrary reference frame in a LI scenario is given by
$\text{d}\Gamma=\frac{1}{8\pi^{2}E_{\bm{p}}}\frac{\text{d}^{3}\bm{p}^{\prime}}{2E_{\bm{p}^{\prime}}}\frac{\text{d}^{3}\bm{k}}{2E_{\bm{k}}}\delta^{(4)}(p-p^{\prime}-k)\overline{|\mathcal{M}(p\to p^{\prime},k)|^{2}}.$ (22)
Here $\bm{p}$ is the $3-$momentum of the initial particle, $\bm{p}^{\prime}$ and $\bm{k}$ are the $3-$momentum of the final particles, and $\mathcal{M}(p\to p^{\prime},k)$ is the Lorentz invariant matrix element which depends on the interaction vertex considered. The Dirac delta function imposes the conservation of the energy and the $3-$momentum, and the $3-$momentum differentials come from the Lorentz invariant phase space (LIPS).
It is important to understand whether equation (22) is allowed to be used in a LI breaking scenario, as is the case with our MDRs. When LI is not broken, the canonical commutation relations between creation and annihilation operators of the field are $[a_{\bm{p}},a_{\bm{p}^{\prime}}^{\dagger}]=(2\pi)^{3}\delta^{(3)}(\bm{p}-\bm{p}^{\prime})$ when the usual factor $(2E_{\bm{p}})^{-1/2}$ in the momentum integral of the field operators is included. This factor on the field operators is reflected in the structure of the LIPS in equation (22). This is no longer true when LI does not hold. Indeed, the wave function of the particles involved in the process will have a different normalization condition, including corrections due to the LI breaking term in the MDR. Hence, to ensure the canonical normalization condition, field operators will no longer be normalized by a factor $(2E_{\bm{p}})^{-1/2}$ and the LIPS in equation (22) should receive additional corrections. Nevertheless, as we shall see during this section, the matrix elements $\mathcal{M}$ that we will consider violate LI, and so no corrections must be taken into account coming from the LIPS as long as we compute the decay rate in the preferred frame. Equation (22) is then valid for the purposes of this work Jacobson_2006 .
Let us now compute the integrals in equation (22). First of all, if we use the conservation of the $3-$momentum we can immediately integrate over $\bm{p}^{\prime}$ and obtain
$\text{d}\Gamma=\frac{1}{8\pi^{2}E_{\bm{p}}}\frac{1}{2E_{\bm{p}-\bm{k}}}\frac{\text{d}^{3}\bm{k}}{2E_{\bm{k}}}\delta(E_{\bm{p}}-E_{\bm{p}-\bm{k}}-E_{\bm{k}})\overline{|\mathcal{M}(p\to p^{\prime},k)|^{2}},$ (23)
where $\mathcal{M}$ is assumed to be evaluated at $p^{\prime}=(E_{\bm{p}-\bm{k}},\bm{p}-\bm{k})$. To integrate over $\bm{k}$ we can use spherical coordinates and fix the $3-$momentum $\bm{p}$ along the $z$ axis. Note that this is possible since the MDR is invariant under rotations (and thus is the decay rate). As will be seen, the matrix element $\mathcal{M}(p\to p^{\prime},k)$ for the processes we consider does not depend on the azimutal angle, so
$\Gamma=\frac{1}{16\pi E_{\bm{p}}}\int\text{d}k\,k^{2}\int_{-1}^{1}\text{d}\cos{\theta}\,\frac{1}{E_{\bm{p}-\bm{k}}E_{\bm{k}}}\delta(E_{\bm{p}}-E_{\bm{p}-\bm{k}}-E_{\bm{k}})\overline{|\mathcal{M}(p\to p^{\prime},k)|^{2}}.$ (24)
On the other hand, elemental properties of the Dirac delta function allow us to write
$\delta(E_{\bm{p}}-E_{\bm{p}-\bm{k}}-E_{\bm{k}})=2E_{\bm{p}-\bm{k}}\delta(E_{\bm{p}-\bm{k}}^{2}-[E_{\bm{p}}-E_{\bm{k}}]^{2}).$ (25)
When deriving the threshold condition in section 2, it can be shown from equation (11) that up to order $\mathcal{O}(\Theta_{c})$
$E_{\bm{p}-\bm{k}}^{2}-[E_{\bm{p}}-E_{\bm{k}}]^{2}=2pk\left(\cos{\theta_{c}}-1+\Theta_{c}\right),$ (26)
which substituting in equation (25) gives
$\delta(E_{\bm{p}}-E_{\bm{p}-\bm{k}}-E_{\bm{k}})=\frac{E_{\bm{p}-\bm{k}}}{pk}\delta(\cos{\theta_{c}}-1+\Theta_{c}).$ ) . (27)
Finally, integrating over $\cos{\theta_{c}}$ ) one gets
$\Gamma=\frac{1}{16\pi p^{2}}\int\text{d}k\,H(\Theta_{c})\overline{|\mathcal{M}(p\to p^{\prime},k)|^{2}}.$ (28)
Note that we have approximated $E_{\bm{p}}=p+\mathcal{O}(\mathbb{A})$ in the denominator for the same reason why we did not consider higher order terms arising from the LIPS. The integration limits for $k$ are determined by the threshold condition (14), being $k=0$ and $k=p$ the lowest and highest integration limits possible, respectively. $H(\Theta_{c})$ is the Heaviside function and establishes whether the process is kinematically allowed or forbidden.
Equation (28) is valid for any process $a\to a+b$ with the MDR proposed with a matrix element $\mathcal{M}$ that does not depend on the azimutal angle. However, the emission rate of the $b$ particles does not allow us to impose any constraint on the LIVPs, unlike the energy loss rate of the $a$ particle. To consider the energy carried off by the particle $b$, we must insert the energy $E_{\bm{k}}=k+\mathcal{O}(\mathbb{B})$ of this particle in the integral of equation (28), obtaining the following energy loss rate of the particle $a$:
$\frac{\text{d}E}{\text{d}t}=\frac{1}{16\pi p^{2}}\int\text{d}k\,kH(\Theta_{c})\overline{|\mathcal{M}(p\to p^{\prime},k)|^{2}}.$ (29)
To evaluate the matrix element $\mathcal{M}$ we will consider, for simplicity, the case where the particle $a$ is a complex scalar field, even though the constraints obtained in section 4 for the LIVPs are applied to fermionic particles. It is expected that spin corrections are of order $\mathcal{O}(1)$ and, therefore, are not significant enough to modify the order of magnitude of the constraints Moore_2001 .
3.2 Electromagnetic Cherenkov radiation
Let us first consider the process described in figure 1 where the particle $b$ is a photon $\gamma$. Feynman rules for a charged complex scalar field applied to this process give the following matrix element schwartz2013quantum :
$i\mathcal{M}=-ie_{a}(p^{\mu}+p^{\prime\mu})\epsilon_{\mu}(k).$ (30)
Here $\epsilon_{\mu}$ is the polarization vector of the photon, and $e_{a}$ the electric charge of the particle $a$. The analysis we will carry out does not distinguish between polarizations, hence we must sum over the two transverse polarization states of the photon when computing the squared matrix element. Using the conservation of $3-$momentum we get
$\overline{|\mathcal{M}(p\to p^{\prime},k)|^{2}}=\sum_{\epsilon}e_{a}^{2}\left|(2p^{\mu}-k^{\mu})\epsilon_{\mu}(k)\right|^{2}.$ (31)
The physical polarizations are perpendicular to the $4-$momentum of the photon, therefore $k^{\mu}\epsilon_{\mu}(k)=0$. The product $p^{\mu}\epsilon_{\mu}(k)$ is easily calculated taking into account that the two polarizations are perpendicular to each other and the angle between $\bm{p}$ and $\bm{k}$ is $\theta_{c}$. This yields
$\overline{|\mathcal{M}(p\to p^{\prime},k)|^{2}}=4e_{a}^{2}p^{2}\sin^{2}{\theta_{c}}=8e_{a}^{2}p^{2}\Theta_{c}+\mathcal{O}(\Theta_{c}^{2}),$ (32)
where we have used equation (12), considered the LI violating terms up to first order, and again denoted $p=|\bm{p}|$. Substituting this matrix element in equation (29), the electromagnetic energy loss rate is then given by
$\frac{\text{d}E}{\text{d}t}=\frac{e_{a}^{2}}{2\pi}\int\text{d}k\,k\Theta_{c}(k)H(\Theta_{c}).$ (33)
The integral has a simple solution:
$\frac{\text{d}E}{\text{d}t}=\frac{e_{a}^{2}}{2\pi}\Bigg{[}$ $\mathbb{A}p^{\alpha+2}\left(1-\frac{k}{p}\right)^{2}\left\{\frac{1}{2}-\frac{1}{\alpha+3}\left(1-\frac{k}{p}\right)^{\alpha+1}\right\}$
$+$ $\mathbb{B}_{\gamma}k^{\beta+2}\left\{\frac{1}{\beta+2}-\frac{1}{\beta+3}\frac{k}{p}\right\}-\frac{m^{2}k^{2}}{4p^{2}}\Bigg{]}_{k_{\text{min}}}^{k_{\text{max}}},$ (34)
where the integration limits $k_{\text{min}}$ and $k_{\text{max}}$ are obtained by solving the threshold condition (14). Also, note that if $\Theta_{c}(k)<0$ for all $k\in[0,p]$ the Heaviside function in equation (33) imposes the energy loss rate to be equal to zero, as Cherenkov emission is kinematically forbidden.
The key feature of this emission rate is that it is of first order in $\Theta_{c}$, and hence in the LI breaking parameters, $\mathbb{A}$ and $\mathbb{B}$, and in $m^{2}/p^{2}$. We have checked that the results obtained here coincide with those in Moore_2001 when considering $\mathbb{A}=0$ and $\beta=0$. Similar MDRs have been used to study vacuum electromagnetic CR Jacobson_2003 ; Jacobson_2006 and it was shown that for $\alpha=\beta\neq 0$ a significant fraction of the energy of the particle $a$ is emitted almost immediately when $a$ has an energy above the threshold of the process. In Sec. 4.3 we will also check with a simple estimation that when $\alpha\neq\beta$ and neither are equal to 00, we obtain the same result. In particular, when $\alpha=\beta=0$, the energy loss rate is suppressed by the difference of the maximum attainable velocities of the particles, $\text{d}E/\text{d}t\sim(c_{a}^{2}-c_{\gamma}^{2})\alpha_{\text{em}}E^{2}$, where $\alpha_{\text{em}}$ is the fine structure constant Coleman_1999 . Thus, a kinematical analysis is enough to establish stringent constraints to LIVPs and it is not necessary to consider propagation effects, i.e., compute the energy loss rate. In this scenario the study of the threshold condition in Sec. 2.3 should be carried out obtaining the threshold momentum of the particle $a$, namely $p_{\text{th}}$. Constraints can then be imposed using high energy astrophysical observations and considering the energy of the particle $a$ detected as the threshold energy of the process. This has been done for $\alpha=\beta\in\mathbb{Z}^{+}$ Coleman_1997 ; Coleman_1999 ; Jacobson_2003 ; Jacobson_2006 but a generalization to $\alpha\neq\beta$ solving the threshold condition (14) for the momentum of the particle $a$ could be performed for small values of the exponents $\alpha$ and $\beta$.
3.3 Gravitational Cherenkov radiation
We now consider the case where the particle $b$ in figure 1 is a graviton $h$. The Feynman rules for the process $a\to a+h$, with $a$ a complex scalar field can be found in reference Han_1999 :
$i\mathcal{M}$ $=$ $-i\sqrt{4\pi G_{\text{N}}}(m^{2}\eta_{\mu\nu}+C_{\mu\nu\rho\sigma}p^{\rho}p^{\prime\sigma})\epsilon^{\mu\nu},$ (35)
$C_{\mu\nu\rho\sigma}$ $=$ $\eta_{\mu\rho}\eta_{\nu\sigma}+\eta_{\mu\sigma}\eta_{\nu\rho}-\eta_{\mu\nu}\eta_{\rho\sigma}.$ (36)
Here $G_{\text{N}}$ is the gravitational constant, and $\epsilon^{\mu\nu}$ is the polarization tensor for the graviton field $h_{\mu\nu}$. This tensor is traceless ($\epsilon_{\mu\nu}\eta^{\mu\nu}=0$) and transverse ($\epsilon_{\mu\nu}k^{\mu}=0$), and can be constructed in terms of the polarization vectors of massive vector bosons Han_1999 . As in the electromagnetic case, we do not distinguish between the polarizations of the graviton and thus we sum over the two physical polarization states. Using equation (36) and substituting in the matrix element (35) one finds
$\overline{|\mathcal{M}(p\to p^{\prime},k)|^{2}}=16\pi G_{\text{N}}\sum_{\epsilon}|p_{\mu}p_{\nu}\epsilon^{\mu\nu}|^{2}=16\pi G_{\text{N}}p^{4}\sin^{4}{\theta_{c}}.$ (37)
Comparing this expression with that corresponding to the electromagnetic CR process, equation (32), it can be noted that for the gravitational case instead of the electric charge $e_{a}^{2}$ we have the gravitational constant times the $3-$momentum squared $G_{\text{N}}p^{2}$, and the emission is reduced for small opening angles, having $\sin^{4}{\theta_{c}}$ rather than $\sin^{2}{\theta_{c}}$. Substituting the matrix element (37) in the energy loss rate (29) and considering the lowest order in $\Theta_{c}$ in equation (12) we obtain
$\frac{\text{d}E}{\text{d}t}=4G_{\text{N}}p^{2}\int\text{d}k\,k\,\Theta_{c}^{2}(k)H(\Theta_{c}).$ (38)
We first notice that the energy loss rate (38) of the vacuum gravitational CR, due to the tensor nature of the gravitational field $h$, is of order $\mathcal{O}(\Theta_{c}^{2})$, whereas the energy loss rate (33) of the electromagnetic CR is of order $\mathcal{O}(\Theta_{c})$ because of the vector nature of the EM field. On the other hand, the electromagnetic coupling constant is much stronger than the gravitational coupling for energies much lower than the Planck scale: the factor $e_{a}^{2}/(2\pi)$ in equation (33) is typically of order $\mathcal{O}(10^{-2})$, while in contrast the factor $G_{\text{N}}p^{2}$ in (38) is of order $\mathcal{O}(10^{-14})$ for the highest energy particles observed in the universe. Therefore, as expected, the energy loss rate due to gravitational CR is much smaller than that caused by the emission of electromagnetic CR. So, for gravitational CR the decay rate of a particle of astrophysical origin can be comparable to its travel time and, therefore, it must be taken into account in order to constrain the LIVPs. Unlike for the electromagnetic CR, a kinematic analysis is not enough to determine an upper bound for $\mathbb{A}$ and $\mathbb{B}$.
We shall also remark that it is possible to have both electromagnetic and gravitational CR emission at the same time. In that scenario, the electromagnetic process is much more efficient, as we have discussed, and it is the dominant source of energy loss. Hence, constraints coming from gravitational CR can be considered as long as the LIVP for the photon $\mathbb{B}_{\gamma}$ is such that the electromagnetic process is always forbidden.
The indefinite integral in equation (38) can be computed in the general case $m\neq 0$ and $\alpha,\beta\geq 0$ in terms of hypergeometric functions, although it is not particularly illuminating. In addition, we also need to calculate the threshold momentum $k_{\text{th}}$ for the integration limits, which cannot be done analytically for arbitrary values of $\alpha$ and $\beta$. Nevertheless, the energy loss rate of the $a$ particle can be obtained for particular values of the parameters. For example, taking $\mathbb{A}=0$, one gets in the massless limit
$\frac{\text{d}E}{\text{d}t}=\frac{2G_{\text{N}}\mathbb{B}^{2}p^{2(\beta+2)}}{(\beta+2)(\beta+1)(2\beta+3)},$ (39)
as it was obtained in reference Kiyota_2015  [^3] This result corresponds to equation (13) of Kiyota_2015 taking $\delta=0$ and identifying our exponent $\beta$ with $\alpha-2$ (do not confuse this exponent with that of the particle $a$) and our parameter $\mathbb{B}$ with $-A/2$. for $k_{\text{min}}=0$ and $k_{\text{max}}=p$. This result reduces to that presented in reference Moore_2001 if we additionally impose $\beta=0$.
4 Constraints on the Lorentz invariance violating parameters
We are now interested in obtaining estimations for the LIVPs $\mathbb{A}$ and $\mathbb{B}$ appearing in $n_{a}$ and $n_{b}$ using the energy loss rate of vacuum gravitational CR (38). We shall use high energy astrophysical observations for this purpose, arguing that UHECRs offer the most stringent bounds. We will consider different values of $\alpha$ and $\beta$ to discuss the corresponding phenomena and illustrate it in figure 2 for $\alpha=2$, $\beta=0$ and $\alpha=1$, $\beta=2$.
4.1 Constraints from the energy loss rate
In section 3 we discussed that the decay rate of a particle of astrophysical origin due to gravitational CR is expected to be of the same order as the time of propagation, which implies that the energy loss rate has to be taken into account in order to estimate the LIVPs of both the particle $a$ and the theory of gravity. Computing the integral in equation (38) and solving the differential equation would give the maximum time travel $t$ possible for a given momentum $p$ in terms of $\mathbb{A}$, $\mathbb{B}$ Moore_2001 ; Kimura_2012 . Thus, constraints on $\mathbb{A}$ and $\mathbb{B}$ can be imposed using high energy astrophysical observations if the distance travelled $ct$ and the momentum $p$ are known.
However, we have argued that it is not possible to obtain the maximum time travel $t$ for arbitrary values of $\alpha$ and $\beta$ in terms of $\mathbb{A}$ and $\mathbb{B}$, as $k_{\text{th}}$ cannot be computed analytically for arbitrary values of $\alpha$ and $\beta$. In this scenario, the condition at which damping from gravitational CR becomes relevant for a particle with energy $p$ travelling for a time $t$ may be estimated as $\text{d}E/\text{d}t\gg p/t$ Kiyota_2015 . Then, as the $a$ particles arrive to the Earth, we establish that values of $\mathbb{A}$ and $\mathbb{B}$ that satisfy $\text{d}E/\text{d}t\gg p/t$ are excluded by gravitational CR, and those that satisfy $\text{d}E/\text{d}t\ll p/t$ are allowed by observation.
To illustrate how constraints can be estimated, let us consider the simple case where $\alpha=\beta=0$. Attending to equations (17) and (18) we see that the process is allowed when $\mathbb{B}\geq\mathbb{A}$. Note that this condition is equivalent to the particle $a$ having a greater phase velocity than the particle $b$, $c_{a}>c_{b}$. Integration of equation (38) in the massless limit is straightforward and yields
$\frac{\text{d}E}{\text{d}t}=\frac{G_{\text{N}}}{3}p^{4}(\mathbb{A}-\mathbb{B})^{2},$ (40)
so damping from CR is not significant if
$0\leq\mathbb{B}-\mathbb{A}\ll\sqrt{\frac{3}{G_{N}tp^{3}}}.$ (41)
Here we see that constraints are more stringent for high energy particles that have travelled a long distance $ct$; note, however, that the energy of the particle plays a more relevant role since the dependence is of the form $(tp^{3})^{-1/2}$. Another interesting feature is that the constraint is only imposed for the difference between the LIVPs, so one of them remains unfixed: as long as this difference is small enough, the values of $\mathbb{A}$ and $\mathbb{B}$ can be arbitrarily large. One should be careful with this last statement as our results have been derived assuming that $|\mathbb{A}p^{\alpha}|$ and $|\mathbb{B}k^{\beta}|$ are very small compared to unity. We have checked that both of these features, the dependence on $p$ and $t$ and the impossibility of fixing both $\mathbb{A}$ and $\mathbb{B}$ simultaneously, are present for different values of $\alpha$ and $\beta$.
It is also worth to mention the case where $\mathbb{A}=0$ and $\beta$ is a positive real number. Performing a similar calculation in the massless limit $m=0$ shows that damping from gravitational CR is not relevant if
$0<\mathbb{B}^{(\beta)}p^{\beta}\ll\frac{\sqrt{(\beta+1)(\beta+2)(2\beta+3)}}{\sqrt{2G_{\text{N}}tp^{3}}},$ (42)
which corresponds to the result derived in reference Kiyota_2015 . The constraint on $\mathbb{B}^{(\beta)}p^{\beta}$ exhibits the same dependence on the energy and the time travel as equation (41), and we see that the value of $\mathbb{B}^{(\beta)}$ decreases for higher values of $\beta$. This has been used to derive much more stringent constraints in the gravitational sector using the absence of vacuum CR for $\beta\gtrsim 0$ rather than using direct detection of GWs, in particular the events GW150914 and GW151226 Yunes_2016 . The same behaviour is observed when the dispersion relation of $a$ particles is modified, that is, when $\mathbb{A}\neq 0$.
4.2 Observational constraints in the gravitational sector
[…]
In order to obtain realistic bounds Kosteleck_2015 , let us consider that the UHECR is a proton of energy $p\sim 10^{11}\,\text{GeV}$. The source of these particles is still unknown Abreu_2022 but it is expected that are produced in active galactic nuclei; the nearest is found at a few Mpc, so we may take $ct\sim 10\,\text{Mpc}$. Here we do not take into account the spin of the proton and neither its inner structure. Regarding the former, we already have mentioned that spin corrections are not expected to affect the order of magnitude of the constraints on $\mathbb{A}_{\text{proton}}\equiv\mathbb{A}$ and $\mathbb{B}_{\text{graviton}}\equiv\mathbb{B}$. Attending to the latter, a detailed analysis of the gravitational CR emitted by a proton would require to consider its partonic structure. For soft emitted gravitons with $k\to 0$, this inner structure might be neglected and the pointlike approximation is valid Jacobson_2006 ; for hard emitted gravitons with $k\lesssim p$ it becomes important and the emission rate of its constituents should be computed. This would substantially increase the number of LIVPs considered as $\mathbb{A}_{\text{quark}}\neq\mathbb{A}_{\text{gluon}}$ in general, making the analysis more involved. This study is out of the scope of this work, but it is worth mentioning that this has been studied with more simple MDRs which consider $\alpha=\beta=0$ Gagnon_2004 ; Elliott_2005 . In short, we will assume that the proton has its own effective MDR, whose structure could be understood in terms of the composition of the MDRs of its constituents.
Computing the threshold momentum $k_{\text{th}}$ numerically for $m=m_{\text{proton}}$, performing the integral in equation (38) and imposing the condition $\text{d}E/\text{d}t\ll p/t$ we have obtained the allowed $\mathbb{A}-\mathbb{B}$ parameter space for $\alpha=2$, $\beta=0$ and $\alpha=1$, $\beta=2$ shown in figure 2. Values of $\mathbb{A}$ and $\mathbb{B}$ that define the dark-grey region permit Cherenkov emission without significant damping, whereas in the light-grey region the process is kinematically allowed but forbidden by the observation because of the damping effect. In both cases we see that no constraints are imposed in the region with $\mathbb{A}>0$ and $\mathbb{B}<0$, and also when $\mathbb{A},\mathbb{B}>0$ and $\mathbb{B}p^{\beta}-\mathbb{A}p^{\alpha}<0$. The latter is due to equation (18), since the threshold momentum is $k_{\text{min}}>p$. When both $\mathbb{A}$ and $\mathbb{B}$ are negative the region where kinematics does no permit Cherenkov emission is different in both cases. Attending to equation (17), the case where $\alpha=2$ and $\beta=0$ establishes that $\mathbb{B}-3\mathbb{A}p^{2}\geq 0$ so that the threshold momentum is $k_{\text{max}}>0$; by contrast, when $\alpha=1$ and $\beta=2$ the maximum momentum exists when $\mathbb{A}\leq 0$.
Figure 2: Constraints on the model for $\alpha=2$, $\beta=0$ (left panel) and $\alpha=1$, $\beta=2$ (right panel) using an UHECR with $p\sim 10^{11}\,\text{GeV}$ and $ct\sim 10\,\text{Mpc}$. The dark-gray region in both figures are the values of $\mathbb{A}_{\text{proton}}\equiv\mathbb{A}$ and $\mathbb{B}_{\text{graviton}}\equiv\mathbb{B}$ that satisfy the condition $\text{d}E/\text{d}t\ll p/t$. There, the damping of the UHECR due to gravitational CR is not significant, and it is possible for the particle to reach the Earth while emitting gravitons. The light-gray region shows that Cherenkov emission is kinematically allowed but does not satisfy $\text{d}E/\text{d}t\ll p/t$; hence, our model excludes those values of the LIVPs. The white region corresponds to $(\mathbb{A},\mathbb{B})$ where CR is not permitted, as $\Theta_{c}<0$ for all possible values of $k$.
The estimated excluded region appears in both cases for $|\mathbb{A}p^{\alpha}|\sim|\mathbb{B}p^{\beta}|\lesssim 10^{-17}$ when $p\sim 10^{11}\,\text{GeV}$ and $ct\sim 10\,\text{Mpc}$. Other positive integer values of $\alpha$ and $\beta$ exhibit constraints of the same order of magnitude for these energies and travelled distances. This allows us to check that LIVPs significantly modify the threshold condition of the process since $|\mathbb{A}p^{\alpha}|\gg(m/p)^{2}$, see equation (6). Hence, the approximations performed in section 2 are consistent and the analyses in the massless limit are accurate.
As we have previously advanced, figure 2 shows that it is possible to have arbitrarily large values of $\mathbb{A}$ and $\mathbb{B}$ where vacuum gravitational CR is allowed by observation as long as its difference is small enough. We have checked the same happens for many other positive integer values of $\alpha$ and $\beta$. This fact can be understood in the following way. When computing the energy loss rate, equation (38), it is necessary to obtain the integration limits from the threshold condition $\Theta_{c}=0$ for some given $\mathbb{A}$ and $\mathbb{B}$. Considering values of $\mathbb{A}$ and $\mathbb{B}$ where the process is allowed for some $k$, an increase in $\mathbb{A}$ and $\mathbb{B}$ will cause an increase in the value of $\Theta_{c}$ and thus in $\text{d}E/\text{d}t$. However, if the difference between $\mathbb{A}$ and $\mathbb{B}$ is small enough, the region of $k$ where the Cherenkov emission is possible will be narrower; this reduces the value of $\text{d}E/\text{d}t$ as the integration interval is smaller, and it can be such that the condition $\text{d}E/\text{d}t\ll p/t$ is still satisfied. Remember that $\Theta_{c}$ was obtained assuming $|\mathbb{A}p^{\alpha}|\sim|\mathbb{B}k^{\beta}|\ll 1$, and thus our model is not predictive for large values of $\mathbb{A},\mathbb{B}$.
In this analysis we notice that when $\beta>0$, given a small value of $\mathbb{A}<0$ we obtain an upper bound for $\mathbb{B}$ but not a lower bound. This is not the case for positive values of $\mathbb{A}$ and $\mathbb{B}$, since fixing one of them constraints the other LIVP; the same happens for negative values of $\mathbb{A}$ and $\mathbb{B}$ when $\beta=0$. Therefore, constraints derived for $\mathbb{A}$ from a different process may not fix $\mathbb{B}$ for $\beta>0$. Note, however, that given a value of $\mathbb{B}$, determined by direct observation of GWs, it is always possible to obtain an upper bound for $\mathbb{A}$. Unfortunately, these measures of the parameter $\mathbb{B}$ are not stringent, so the bounds obtained for $\mathbb{A}$ are not competitive with those obtained from threshold constraints in the QED sector Jacobson_2003 ; Jacobson_2006 .
4.3 Travel distance for particles emitting electromagnetic Cherenkov radiation
We have previously discussed that for positive integer values of $\alpha=\beta$ the energy loss rate of the particle $a$ due to vacuum electromagnetic CR is large enough so that the travelled distance for energies above the threshold $p_{\text{th}}$ is negligible compared to astrophysical distances Coleman_1997 ; Coleman_1999 ; Jacobson_2003 ; Jacobson_2006 . As our MDR does not assume equal values of the exponents of the particle $a$ and the photon ($\alpha$ and $\beta_{\gamma}$, respectively), it could be possible for $a$ to emit CR over large distances. Here we show that, even when $\alpha\neq\beta$, the vacuum electromagnetic CR is much more efficient than its gravitational counterpart, and that the particle $a$ losses all of its energy almost immediately above the threshold.
Figure 3: Constraints on the model for $\alpha=2$, $\beta=0$ (left panel) and $\alpha=1$, $\beta_{\gamma}=2$ (right panel) using an UHECR with $p\sim 10^{11}\,\text{GeV}$ and $ct\sim 0.1\,\text{fm}$. White, light-gray and dark-gray regions represent the same as in figure 2 but for the existence or absence of vacuum electromagnetic CR.
To illustrate this, let us consider an ultra high energy proton of energy $p\sim 10^{11}\,\text{GeV}$. As we do not solve the threshold equation (14) to find the threshold momentum $p_{\text{th}}$ of the particle $a$, we estimate that the order of magnitude of $\mathbb{A}$ and $\mathbb{B}_{\gamma}$ given by the kinematical constraints is the same as the one obtained in the gravitational case, see figure 2. For the same values of the exponents $\alpha$ and $\beta_{\gamma}$, figure 3 shows the parameter space $\mathbb{A}-\mathbb{B}_{\gamma}$ assuming a travelled distance of $ct\sim 0.1\,\text{fm}$. We see that the region where the damping is not significant lies in distance scales $39$ orders of magnitude smaller compared to those assumed for the emission of vacuum gravitational CR. This means that for UHECRs whose origin is galactic or extragalactic, the emission of electromagnetic CR occurs almost immediately after being produced, losing energy until reaching the threshold momentum $p_{\text{th}}$: the dark-grey region occupies an insignificant amount of the parameter space, and thus the kinematical analysis is enough to obtain bounds on the LIVPs. The estimation was made for two specific pair of values $(\alpha,\beta)$ but we have checked explicitly that similar results are obtained for many other positive integer values.
5 Conclusions
Vacuum Cherenkov emission occurs as an emission process of a massless particle and it is possible when LI is broken Schreck_2018 ; Moore_2001 ; Mattingly_2005 ; Coleman_1997 ; Coleman_1999 ; Jacobson_2003 ; Jacobson_2006 ; Gagnon_2004 ; Elliott_2005 ; Kimura_2012 ; Kiyota_2015 ; Kosteleck_2015 . An efficient way to implement LIV is through MDRs, which are able to encapsulate QG phenomenology Addazi_2022 ; batista2023white . In this work we have proposed a new class of LI violating MDRs inspired on the classical electromagnetic Cherenkov emission in an optical medium with refractive index $n$, different for each particle species and with a power law dependence on the energy. Kinematics of the GCR process have been studied and the momentum configurations for which the process is allowed have been derived. These conditions were implemented to obtain the energy loss rate due to GCR, paying particular attention to the electromagnetic and gravitational vacuum Cherenkov radiation. The latter results served to estimate the LIVPs of UHECRs.
We would like to remark an important feature: the parameter space of the LI violating terms is not bounded, which implies that gravitational vacuum Cherenkov radiation is permitted as long as the difference between them is small enough. Different processes, such as the decay of a graviton in two high energy particles, could be used to restrict the allowed parameter space, but the quantum nature of gravity has not been resolved yet and no observations of gravitons have been made. Note that this problem is not present in the QED sector. On the one hand, the characteristic time of the analog processes mentioned is much smaller than the travelled distance of astrophysical particles; hence, constraints can be imposed using only kinematics. On the other hand, electrons and photons of high energy have been detected, and thus the parameter space of the LIVPs of these particles can be restricted from the absence of the vacuum electromagnetic CR and photon decay processes, or the change in the threshold of the photon annihilation process Jacobson_2003 ; Jacobson_2006 .
It should be noted that even though the present work is focused on the study of threshold effects and more specifically particle decays, MDRs have a much broader scope to restrict LIVPs. Constraints in the gravitational sector may be imposed using time of flight delay in multi-messenger detections. At the present only one multi-messenger detection mixing GWs and electromagnetic radiation has been reported, the GW signal GW170817 with the gamma-ray signal from a kilonova Abbott_2017 , but this restrictions using the time delay between GWs and photons have been applied only to impose limits on the Standard Model extension parameters in reference Kosteleck_2015 . Another possibility, which is not considered in the MDR proposed here, is the vacuum birefringence phenomenon caused by the difference in the propagation velocity between the states $h_{+}$ and $h_{\times}$ of GWs.
In conclusion, our work establishes a systematic method to study two body particle decays in a LI violating scenario and impose constraints through high energy astrophysical observations. Extensions to more particle interaction processes can be carried out following the steps exemplified, but these are not of much interest in the gravitational sector. We also want to emphasize the importance of using complementary observations, such as direct detection of GWs, in order to improve the bounds on the parameters introduced. This will be reinforced with future detections of UHECRs and multi-messenger events, which may allow us to understand the elusive nature of QG.
Acknowledgements.
P.M.M. would like to thank Iarley P. Lobo for pointing out to her interesting references. The research of P.M.M. is supported by the project PID2022-138263NB-I00 funded by MICIN/AEI/10.13039/ 501100011033 and by ERDF/EU. The work of J.A.R.C. is partially supported by the project PID2022-139841NB-I00 funded by MICIU/AEI/10.13039/501100011033 and by ERDF/EU, and by COST (European Cooperation in Science and Technology) Actions CA22113 and CA21136. M.A. acknowledges financial support through IPARCOS Master grant.